OpenStudy (anonymous):

3x-y=11 2x+5y=13 how to solve using addition?

jimthompson5910 (jim_thompson5910):

Multiply every term in the first equation by 5 to get 15x-5y=55 Then add this new equation to the second equation 15x-5y=55 2x+5y=13 ------------ 17x+0y=68 So we get 17x = 68

jimthompson5910 (jim_thompson5910):

Solve 17x = 68 for x Then use this solution to find y

OpenStudy (anonymous):

so do i do 17x =68 then i divide then i get x=4 then i plug in the 4?

jimthompson5910 (jim_thompson5910):

Yes, x = 4 Now that you know that x = 4, you can plug it into any equation that has a 'y' in it. For instance, plug it into the second equation 2x+5y = 13 to get 2(4) + 5y = 13 Now solve for y

OpenStudy (anonymous):

y=1?

jimthompson5910 (jim_thompson5910):

You got it